# Improving absolute separability bounds for arbitrary dimensions

### v1

@@ 4 Multipartite absolute separability
In this section, we extend some of the previous results to the multipartite setting. A state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M}\otimes\dots\otimes\mathbb{C}^{T})$ is (fully) separable if it admits a convex decomposition onto product states
$\rho=\sum_{i}p_{i}\outerproduct{e_{i},f_{i},\dots,g_{i}}{e_{i},f_{i},\dots,g_{i}},$ (10)
@@
$\sigma_{N-k}\hskip 4.33382pt\geq\hskip 4.33382pt\frac{d^{k}\cdot\mathds{1}}{d^{N}+2}\hskip 4.33382pt\geq\hskip 4.33382pt\frac{\mathds{1}}{d^{N-k}+2}.$ (11)
However, we recall that certifying separability with respect to all bipartitions in the state and its successive reduced matrices does not imply full separability [38, 39]. That is to say, certain multipartite states are entangled despite being separable with respect to its bipartitions. A paradigmatic example of such states is constructed from unextendible product bases (UPB) [40]. Biseparable entangled states also exist and have been characterized in the symmetric sector [41, 42].
The previous observation evidences the challenges to certify full separability as a much more stringent condition than biseparability. We begin by extending the concept of the Gurvits-Barnum ball [Eq. (1)] to the multipartite setting [43]. This norm-based condition represents one of the few known sufficient criteria for full separability, as we discuss below.
@@
For the specific case of $N$ qubits (i.e., $d=2$), this bound can be improved to $A=\frac{\beta\cdot 2^{N}}{\beta+3^{N}}$, where $\beta=54/17$ [17].
In the same spirit as Theorem 3, Lemma 5 allows us to formalize the following sufficient condition for AS, stemming from the reduction-like map Eq. (2).
@@
then $\sigma=\Lambda_{\alpha}(\rho)$ is fully separable, where $\alpha_{\pm}^{*}=\frac{A\pm\sqrt{A\cdot(D-1)\cdot(D-A)}}{D-A-1}$ and $A=2^{2-N}$.
For the specific case $d=2$ we have $A=\frac{\beta\cdot 2^{N}}{\beta+3^{N}}$, being $\beta=54/17$.
@@
$\lambda_{\min}(\sigma)\geq\frac{1}{d^{N}+\alpha_{+}}\mbox{ or }\lambda_{\max}(\sigma)\leq\frac{1}{d^{N}-\alpha_{-}},$ (14)
where $\alpha_{\pm}$ are defined in Theorem 6, then $\sigma$ is fully separable.
@@
The proof in analogous to that of Theorem 3. ∎
@@
The existence of multipartite states fulfilling Theorem 3 but remaining unconsidered by Lemma 5 introduces the notion of a subtle form of entanglement: states that are AS and AP with respect to any bipartition and with respect to any bipartition of any of the reduced matrices, but yet not fully separable. In light of the intricacies of multipartite AS, in the next section we focus on the symmetric subspace, in which the characterization of these features is more manageable.
@@ 4.1 Symmetric absolute separability and absolute PPT
The notions of AS and AP introduced earlier may be immediately restricted to the symmetric subspace, which we shall denote here as $S(\mathcal{H})$. In a generic $N-$qudit system, $\mathcal{H}=(\mathbb{C}^{d})^{\otimes N}$, this subspace has dimension $\binom{N+d-1}{d-1}$. Specifically, we define
@@
A symmetric state $\rho$ acting on $S(\mathcal{H})$ is symmetric absolutely separable (or SAS) if and only if, given any unitary matrix acting on the symmetric subspace $U_{S}$, $\rho^{\prime}=U_{S}\rho_{S}U_{S}^{{\dagger}}$ is (fully) separable.
In addition, we introduce the weaker concept of symmetric absolutely PPT
@@
A symmetric state $\rho$ acting on $S(\mathcal{H})$ is symmetric absolutely PPT (or SAP) if and only if, given any unitary matrix acting on the symmetric subspace $U_{S}$, $\rho^{\prime}=U_{S}\rho_{S}U_{S}^{{\dagger}}$ is PPT with respect to any bipartition.
Let us start with the simplest case corresponding to the symmetric subspace of a $2-$qubit system. Following the approach from Theorem 2, that is, considering the inverse of the reduction map (3), the corresponding bounds on $\alpha\in[-3/4,1]$ that allow to certify SAS have been already derived in the general context of separability [11]. For SAS, such bounds can be reformulated as:
@@
Let $\rho\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a $2-$qubit system with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{4}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{4}{9},$ (15)
then, $\rho$ is symmetric absolutely separable.
@@
The proof is analogous to that of Theorem 3, and uses the derivations on the inverse of the reduction map introduced by some of us in [11]. ∎
From the two simplexes described by the minimal and maximal eigenvalue conditions, we derive the following sufficient condition for SAS, as done similarly in Proposition 1.
@@
Let $\rho\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a $2-$qubit system with $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}$ its corresponding eigenvalues in non-decreasing order. If
$7\lambda_{0}+5\lambda_{1}\geq 3,$ (16)
then $\rho$ is symmetric absolutely separable.
Proof.
In Appendix A we show how to compute the linear inequality and verify how it is the most external one enclosing the ordered CH under ordered eigenvalues. ∎
However, as appreciated in Figure 3, even though the simplexes are tight, the CH of both of them does not attain the criterion $\sqrt{\lambda_{1}}+\sqrt{\lambda_{0}}\geq 1$ (see [19, 44]), which fully characterizes SAS in the simple $2-$qubit scenario.
Figure 3: Left panel: Comparison of the states detected for a symmetric $2-$qubit system with the convex hull (green line) of the sufficient conditions through the inverse of the linear map with $\alpha=1$ (dashed red line) and $\alpha=-3/4$ (dotted orange line) with the complete characterization $\sqrt{\lambda_{0}}+\sqrt{\lambda_{1}}\geq 0$ (black solid line). Right panel: Comparison of the states detected for a symmetric $3-$qubit system with the convex hull (green line) of the sufficient conditions through the inverse of the linear map with $\alpha=2/3$ (dashed red line) and $\alpha=-2/3$ (dotted orange line) with the best known criterion [45] (black sphere). The representation shows a $2D$ cut of the non-increasingly ordered region of eigenvalues $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$.
Next, for the second-simplest case of a symmetric $3-$qubit system, the bounds of $\alpha\in[-1/3,2/3]$ have also been derived analytically [11], allowing us to prove SAS from the inverse of the reduction map $\Lambda_{\alpha}^{-1}(\rho)$ (3). Nevertheless, numerical evidence [44] shows that the state with eigenvalues $(1,3,3,3)/10=(1-(N)/(N+1+\alpha),1/(N+1+\alpha),1/(N+1+\alpha),1/(N+1+\alpha))$ is still SAS. From this discussion, we infer that the tighter lower bound $\alpha=-2/3$ can be used to prove SAS.
Theorem 8.
Let $\rho\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a $3-$qubit system with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{3}{14}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{3}{10},$ (17)
then, $\rho$ is symmetric absolutely separable.
Proof.
The proof is analogous to the one in Theorem 3. The positive bound of $\alpha=2/3$ has been derived in Ref [11], whereas the negative bound of $\alpha=-2/3$ has been obtained numerically. It is enough to prove it for pure states $\rho=\outerproduct{\Psi}{\Psi}$ |. For $3-$qubits, |$\ket{\Psi}$ has $4$ complex coefficients in the symmetric (Dicke) basis |$\ket{\Psi}=\sum_{i}x_{i}\ket{D_{i}^{(3)}}$. It is possible to minimize for all $\alpha\in[-1,0]$ over the coefficients $\{x_{i}\}$ the smallest eigenvalue of $\Lambda_{\alpha}^{T_{2:4}}(\rho)$ and check that $\alpha=-2/3$ is the smallest value of $\alpha$ that cannot lead to a negative partial transpose. The value is also considered extreme in [44], and we have extended their numerical calculations of the maximal possible negativity for all the states with eigenvalues $(1,3,3,3)/10$ with over 2 million random unitary matrices. ∎
Remarkably, by characterizing the CH of both simplexes, we provide a stronger sufficient condition for SAS from few eigenvalues.
Proposition 4.
Let $\rho\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a $3-$qubit system and $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$ its corresponding eigenvalues in non-decreasing order. If
$6\lambda_{0}+6\lambda_{1}+2\lambda_{2}\geq 3,$ (18)
then $\rho$ is symmetric absolutely separable.
Proof.
The proof is analogous to the one of Proposition 3. For the analytical lower bound of $\alpha\in[-1/3,2/3]$ derived in [11], the linear CH obtained is given by the expression $5\cdot\lambda_{1}+9\cdot\lambda_{0}\geq 3$. ∎
Interestingly, for $N=3$ symmetric qubits, some AS states lie outside the region given by Eq. (18), as recently reported [44]. Nevertheless, our analytical linear expression includes all known, to date, analytical criteria [45], and provides the best existing bound for SAS for this specific set of states. In Figure 3, we display the known conditions that detect SAS states, together with those derived in this work. As it can be seen, the volume of SAS-states detected with our conditions represents a notable improvement.
Regarding SAP (symmetric absolute PPT), let us recall that in the symmetric subspaces $S(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ for $d=2,3$, PPT is a sufficient condition for separability, and thus SAS is equivalent to SAP in such cases. In what follows, we derive new bounds to certify SAP for the general $N-$qudit symmetric spaces with respect to any bipartitation. The specific case of SAP for bipartite qudit systems, i.e. $N=2$, has been completely characterized by a series of matrix inequalities on the eigenvalues of such systems [19]. However, the number of such matrices grows fast with the local dimension $d$. Thus, the conditions we present here are much more practical.
Theorem 9.
[…]
@@
$-\binom{N}{\lfloor N/2\rfloor}^{-1}=-\lambda_{min}\leq\alpha\leq 2\lambda_{min}=2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1},$
where $\lambda_{min}>0$ is the minimal eigenvalue of $\mathds{1}_{S}^{T_{A}}$ and the partition $A$ includes $\lfloor N/2\rfloor$ qudits. Then, $\rho_{S}\geq 0\rightarrow\Lambda_{\alpha}(\rho_{S})=\sigma_{S}$ is SAP.
Similarly, if $\sigma_{S}\geq 0$ and $\Lambda_{\alpha}^{-1}(\sigma_{S})=\rho_{S}\geq 0$, then $\sigma_{S}$ is SAP.
@@
Some interesting remarks stem from Theorem 9. First, symmetric states of $N-$qudits that are absolutely PPT in the partition $11:N-$ 1 can still be NPT with respect to some larger partition. Since the set of SAP states with respect to the largest bipartition is the smallest one, and it is contained in the sets of SAP of smaller partitions, we will consider it as the one defining SAP 4. It is also necessary to mention that we have numerical evidence that the upper bound of $\alpha_{+}=2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1}$ is tight for the SAP set, since increasing such value by even a $0.1\%$ leads to NPT states under certain unitary transformations. Nevertheless, the same behavior has not been found for the negative lower bound on $\alpha_{-}$. As an example, one can check that for $N=2,3$ qubits, the lowest bounds are given by $\alpha_{-}=-3/4,-2/3$ respectively, instead of the values $\alpha_{-}=-1/2,-1/3$ given by Theorem 9.
The previous theorem leads to the following corollary as a sufficient condition for SAP in $N-$qudit systems.
Corollary 2.
Let $\rho\in\mathcal{S}((\mathbb{C}^{d})^{\otimes N})$ be a symmetric state of an $N-$qudit system of dimension $D=d^{N}$ with minimal and maximal eigenvalue $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{d^{N}+2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1}}\text{ or }\lambda_{\max}(\rho)\leq\frac{1}{d^{N}-\binom{N}{\lfloor N/2\rfloor}^{-1}},$ (19)
then $\rho$ is symmetric absolutely PPT.
@@
The proof is analogous to that of Theorem 3. ∎
For the specific case of N-qubit systems, a general linear condition on the eigenvalues of a given state $\rho$ can be provided to ensure it is SAP, in a complete analogy as it was demonstrated in Proposition 1. Moreover, the same procedure can be generalized for N- qudit systems, $S((\mathbb{C}^{d}))^{\otimes N})$, even though we do not explicitly describe it in this work.
Proposition 5.
Let $\rho\in\mathcal{S}((\mathbb{C}^{2})^{\otimes N})$ be a symmetric state of $N$ qubits and $\{\lambda_{i}\}_{i=0}^{N}$ its eigenvalues in non-decreasing order $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,1,..,N}$. If $\rho$ is contained in the CH of the conditions presented in Corollary 2, then
$3\cdot\binom{N}{\lfloor N/2\rfloor}\sum_{i=0}^{\lfloor\frac{N+1}{3}\rfloor-1}\lambda_{i}+\left[\binom{N}{\lfloor N/2\rfloor}\cdot(N+1-3\cdot\lfloor\frac{N+1}{3}\rfloor)+2\right]\cdot\lambda_{\lfloor\frac{N+1}{3}\rfloor}\geq\binom{N}{\lfloor N/2\rfloor}.$ (20)

### v2

@@ 4 Multipartite absolute separability
For multipartite settings, much less is known concerning (absolute) separability. We recall that a state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M}\otimes\dots\otimes\mathbb{C}^{T})$ is (fully) separable if it admits a convex decomposition onto product states
$|,\rho=\sum_{i}p_{i}\outerproduct{e_{i},f_{i},\dots,g_{i}}{e_{i},f_{i},\dots,g_{i}}$, (13)
@@
[…]
$\sigma_{N-k}\hskip 3.4481pt\geq\hskip 3.4481pt\frac{d^{k}\cdot\mathds{1}}{d^{N}+2}\hskip 3.4481pt\geq\hskip 3.4481pt\frac{\mathds{1}}{d^{N-k}+2}.$ (14)
However, we recall that certifying separability with respect to all bipartitions in the state and its successive reduced matrices does not imply full separability [43, 44]. That is to say, certain multipartite states are entangled despite being separable with respect to its bipartitions. Paradigmatic examples of such states can be built from unextendible product bases (UPB) [45] or are given by multiqubit PPT entangled states in the symmetric sector [46, 47].
The previous observation evidences the challenges to certify full separability as a much more stringent condition than biseparability. A bound on the exponent of the radius of the multipartite fully separable ball (Gurvits-Barnum) was provided [48]. This norm-based condition represents one of the few known sufficient criteria for full absolute separability.
@@
For the specific case of qubits, $d=2$, this bound can be improved to $A=\frac{\beta\cdot 2^{N}}{\beta+3^{N}}$, where $\beta=54/17$ [20]. In the same spirit as Theorem 3, Lemma 2 allows us to formalize the following sufficient condition for AS, stemming from the reduction-like map Eq. (2).
@@
then $\sigma=\Lambda_{\alpha}(\rho)$ is absolutely fully separable, where $\alpha_{\pm}^{*}=\frac{A\pm\sqrt{A\cdot(D-1)\cdot(D-A)}}{D-A-1}$ and $A=2^{2-N}$.
For the specific case $d=2$ we have $A=\frac{\beta\cdot 2^{N}}{\beta+3^{N}}$, where $\beta=54/17$.
@@
$\lambda_{\min}(\sigma)\geq\frac{1}{d^{N}+\alpha_{+}}\quad\text{or}\quad\lambda_{\max}(\sigma)\leq\frac{1}{d^{N}+\alpha_{-}},$ (17)
where $\alpha_{\pm}$ are defined in Theorem 6, then $\sigma$ is absolutely fully separable.
@@
The proof is analogous to that of Theorem 3. ∎
@@
The existence of multipartite states fulfilling Theorem 3 but remaining unconsidered by Lemma 2 introduces the notion of a subtle form of entanglement: states that are AS and AP with respect to any bipartition and with respect to any bipartition of any of the reduced matrices, but yet not fully separable. In light of the intricacies of multipartite AS, in the next section we focus on the symmetric subspace, in which the characterization of these features is more manageable, as symmetric states are either fully entangled or fully separable.
@@ 4.1 Symmetric absolute separability and absolute PPT
The notions of AS and AP introduced earlier may be immediately restricted to the symmetric subspace, which we shall denote here as $S(\mathcal{H})$. In a generic NN-qudit system, $\mathcal{H}=(\mathbb{C}^{d})^{\otimes N}$, this subspace has dimension $\binom{N+d-1}{d-1}$. Specifically, we define
@@
A symmetric state $\rho$ acting on $S(\mathcal{H})$ is symmetric absolutely separable (SAS) if and only if, given any unitary matrix acting on the symmetric subspace $U_{S}$, $\rho^{\prime}=U_{S}\rho_{S}U_{S}^{{\dagger}}$ is fully separable.
@@
A symmetric state $\rho$ acting on $S(\mathcal{H})$ is symmetric absolutely PPT (SAP) if and only if, given any unitary matrix acting on the symmetric subspace $U_{S}$, $\rho^{\prime}=U_{S}\rho_{S}U_{S}^{{\dagger}}$ is PPT with respect to any bipartition.
The symmetric subspace is of particular significance for the application of the techniques presented before to certify a given convex property from the spectrum of the state. In what follows, we provide sufficient SAS and SAP criteria by studying the tight bounds of $\alpha_{\pm}$ from the inverse map condition Eq. (3) and the convexity criteria presented in Lemma 1. For this case, it is convenient to work in the lower-dimensional Dicke basis [49] directly with symmetric matrices $\rho_{S}$ around the symmetric projector $\mathds{1}_{S}.$
First, we will consider the symmetric cases of $2$- and $3$-qubit systems. Even though they are simple cases, we already provide a better analytical characterization for the latter case than what is known. Following the approach from Theorem 2, that is, considering the inverse of the reduction map (3) in the symmetric subspace, the $2$-qubit bounds on $\alpha\in[-3/4,1]$ that allow to certify SAS have been already derived in the general context of separability [12]. Then, the different SAS criteria can be reformulated as in the following Theorem 7.
@@
Let $\rho_{S}\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a 22-qubit system with $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}$ its corresponding eigenvalues in non-decreasing order. If
$\lambda_{0}\geq\frac{1}{4}\quad\text{or}\quad\lambda_{2}\leq\frac{4}{9},$ (18)
then $\rho$ is SAS.
Moreover, if
$7\lambda_{0}+5\lambda_{1}\geq 3,$ (19)
then $\rho$ is SAS.
@@
The proof of the single eigenvalue expressions is analogous to that of Theorem 3, and uses the derivations on the inverse of the reduction map introduced by some of us in [12]. The linear inequality follows from Lemma 1 by considering $\alpha_{+}=1$, $\alpha_{-}=-3/4$ and $D=3$. ∎
However, as appreciated in Figure 3, even though the simplexes are tight, the CH of both of them does not attain the criterion $\sqrt{\lambda_{1}}+\sqrt{\lambda_{0}}\geq 1$ (see [22, 50]), which fully characterizes SAS in the simple symmetric 22-qubit scenario. For the case of a symmetric 33-qubit system, we managed to provide the actual tight bounds of $\alpha_{\pm}$, which had also been obtained numerically in [50].
Figure 3: Left panel: Comparison of the states detected for a symmetric 22-qubit system with the convex hull (green line) of the sufficient conditions through the inverse of the linear map with $\alpha=1$ (dashed red line) and $\alpha=-3/4$ (dotted orange line) with the complete characterization $\sqrt{\lambda_{0}}+\sqrt{\lambda_{1}}\geq 1$ (black solid line). Right panel: Comparison of the states detected for a symmetric 33-qubit system with the convex hull (green line) of the sufficient conditions through the inverse of the linear map with $\alpha=2/3$ (dashed red line) and $\alpha=-2/3$ (dotted orange line) with the best known analytical criterion [25] (black sphere). The representation shows a $2D$ cut of the non-increasingly ordered region of eigenvalues $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$.
Theorem 8.
Let $\rho_{S}\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a 33-qubit system and $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$ its corresponding eigenvalues in non-decreasing order. If
$\lambda_{0}\geq\frac{3}{10}\quad\text{or}\quad\lambda_{3}\leq\frac{3}{14},$ (20)
then $\rho$ is AS.
Moreover, if
$6\lambda_{0}+6\lambda_{1}+2\lambda_{2}\geq 3,$ (21)
then $\rho$ is SAS.
Proof.
Eq. (20) is analogous to the ones in Theorem 3. For the positive bound of $\alpha=2/3$ see Proposition 2 , and for $\alpha_{-}=-2/3$ see Theorem 10. The proof of Eq. (21) follows from Lemma 1 by considering the announced $\alpha_{\pm}$ and $D=4$. ∎
Interestingly, for $N=3$ symmetric qubits, some SAS states, found numerically, lie outside the region given by Eq. (21), as recently reported in [50]. Nevertheless, our analytical linear expression includes all known, to date, analytical criteria [25, 51]. In Figure 3, we display the known conditions that detect SAS states, together with those derived in this work. As it can be seen, the volume of SAS states detected with our conditions represents a notable improvement.
For higher dimensional cases, we will focus on the SAP (symmetric absolute PPT) set of states. Let us recall that in the symmetric subspaces $S((\mathbb{C}^{2})^{\otimes N})$ for $N=2,3$, PPT is a sufficient condition for separability, and thus SAS is equivalent to SAP in such cases. In what follows, we derive new bounds to certify SAP for the general NN-qudit symmetric spaces with respect to any bipartitation. Very recently, some counterexamples for the presumed equivalence between the SAS and SAP sets of states have been found numerically for $N=5,7,9$ in qubit systems [52]. Nevertheless, the SAP set of states has only been characterized for the specific case of bipartite qudit systems, i.e., $N=2$, via a set of LMI on the eigenvalues of such systems [22]. However, the number of such inequalities grows with the local dimension $d$ as $(d(d+1)/2)!$ becoming intractable already for $d=5$. Here, we present new conditions valid for arbitrary dimensions and number of parties $d,N$, which are much more practical.
Regarding SAP, it is also necessary to address that absolute positive partial transposition in multipartite systems is defined w.r.t. all the bipartitions of the state at once. For symmetric multipartite states, it can be seen (c.f. Proposition 2 ) that the sets of SAP fulfill the inclusions $\text{SAP}_{\lfloor N/2\rfloor:N-\lfloor N/2\rfloor}\in\text{SAP}_{\lfloor N/2\rfloor-1:N-\lfloor N/2\rfloor+1}\in\cdots\in\text{SAP}_{2:N-2}\in\text{SAP}_{1:N-1}$. Thus, it is necessary to consider SAP w.r.t. the biggest (or most balanced) bipartition , otherwise it is always possible to find a state that is PPT under the action of any symmetric unitary in the biggest bipartition and NPT for some symmetric unitary in the smaller one. To this aim, in the following Proposition 2 , we derive analytical bounds for $\alpha_{\pm}$ to certify SAP in $N$-qudit systems.
@@
Let $\rho_{S}\in\mathcal{S}((\mathbb{C}^{d})^{\otimes N})$ be a symmetric state of an $N$-qudit system with $\{\lambda_{k}\leq\lambda_{k+1}\}_{k=0,\cdots,D_{S}-1}$ its corresponding eigenvalues in non-decreasing order, with $D_{S}=\binom{N+d-1}{d-1}$. If at least one of the following inequalities holds:
@@
$\lambda_{0}\geq\frac{1}{D_{S}+2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1}}\quad\text{or}\quad\lambda_{D_{S}-1}\leq\frac{1}{D_{S}-\binom{N}{\lfloor N/2\rfloor}^{-1}},$ (22)
then $\rho_{S}$ is symmetric absolutely PPT.
Moreover, if
$3\binom{N}{\lfloor N/2\rfloor}\sum_{i=0}^{\left\lceil\frac{D_{S}-2}{3}\right\rceil-1}\lambda_{i}+\left[\binom{N}{\lfloor N/2\rfloor}\left(D_{S}-3\left\lceil\frac{D_{S}-2}{3}\right\rceil\right)+2\right]\lambda_{\left\lceil\frac{D_{S}-2}{3}\right\rceil}\geq\binom{N}{\lfloor N/2\rfloor}$ (23)
then $\rho_{S}$ is symmetric absolutely PPT.
@@
On the other hand, the proof of Eq. (23) follows from Lemma 1 by considering the anounced $\alpha_{+}=2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1}$, $\alpha_{-}=-\binom{N}{\lfloor N/2\rfloor}^{-1}$ and $D=D_{S}=\binom{N+d-1}{d-1}$. We also considered the simplification $c=\left\lceil\frac{1}{3}\left(-3+2\binom{N}{\lfloor N/2\rfloor}^{-1}+D_{S}\right)\right\rceil=\left\lceil\frac{D_{S}-2}{3}\right\rceil$. ∎
An interesting remark is that we have numerical evidence that the upper bound of $\alpha_{+}=2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1}$ is tight for the SAP set, since increasing such value by even a $0.1\%$ leads to NPT states under certain unitary transformations. Nevertheless, the same behavior has not been found for the negative lower bound on $\alpha_{-}=\binom{N}{\lfloor N/2\rfloor}^{-1}$, we foresee a dependence on $d$ for $\alpha_{-}$. Thus, the states that lead to a partial transpose with maximal eigenvalue $1$ do not fall in the span of the eigenvectors of $\mathds{1}_{S}$ with minimal eigenvalue [53, 54]. As an example, one can check that for $N=2,3$ qubits, the lowest bounds are given by $\alpha_{-}=-3/4,-2/3$ respectively, instead of the values $\alpha_{-}=-1/2,-1/3$ given by Proposition 2 . For a given $N$, the negative bound becomes tighter as $d\rightarrow\infty$. Nevertheless, in a case by case basis, it is possible to study the structure of the
$\Lambda_{\alpha}(\rho_{S})^{T_{\lfloor N/2\rfloor:N-\lfloor N/2\rfloor}}=\mathds{1}_{S}^{T_{\lfloor N/2\rfloor:N-\lfloor N/2\rfloor}}+\alpha\rho_{S}^{T_{\lfloor N/2\rfloor:N-\lfloor N/2\rfloor}}$ (24)
matrix to ensure positivity and derive better bounds of $\alpha_{-}$ as we present next.
Theorem 9.
Let $\rho_{S}\in\mathcal{S}(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ be a symmetric state of a $2$-qudit system with $\{\lambda_{k}\leq\lambda_{k+1}\}_{k=0,\cdots,D_{S}-1}$ its corresponding eigenvalues in non-decreasing order, with $D_{S}=d\cdot(d+1)/2$. If at least one of the following inequalities holds:
$\lambda_{0}\geq\frac{1}{D_{S}+1}\quad\text{or}\quad\lambda_{D_{S}-1}\leq\frac{1}{D_{S}-\frac{d+1}{2d}},$ (25)
then $\rho_{S}$ is symmetric absolutely PPT.
Moreover, if
$\frac{3d+1}{d+1}\cdot\sum_{i=0}^{c-1}\lambda_{i}+\left[\frac{d(d+1)}{2}-\frac{3d+1}{d+1}\cdot c+1\right]\cdot\lambda_{c}\geq 1$ (26)
where $c=\left\lceil\frac{d\cdot(d+3)\cdot(d-1)}{2\cdot(3d+1)}\right\rceil$, then $\rho_{S}$ is symmetric absolutely PPT.
@@
[…]
$\sigma^{T_{A}}=\bigoplus_{i>j}\begin{pmatrix}\frac{1}{2}&\alpha x_{i}x_{j}\\ \alpha x_{i}x_{j}&\frac{1}{2}\end{pmatrix}\oplus\begin{pmatrix}1+\alpha x_{0}^{2}&\frac{1}{2}&\cdots&\frac{1}{2}\\ \frac{1}{2}&\ddots&&\frac{1}{2}\\ \vdots&&\ddots&\vdots\\ \frac{1}{2}&\frac{1}{2}&\cdots&1+\alpha x_{d-1}^{2}\end{pmatrix}$. (27)
Positivity of the first d(d1)/2d(d-1)/2 blocks implies $1/2-|\alpha|x_{d-1}x_{d-2}\geq 0$. From $x_{d-1}x_{d-2}\leq 1/2$, it yields the range $1\leq\alpha\leq 1$. On the other hand, by convexity, it is sufficient to set $x_{d-2}=0,x_{d-1}=1$ to check the positivity of the last $d\times d$ block. For $\alpha\geq 0$, the block is PSD. The minimal value of $\alpha$, $\alpha_{\min}<0$ such that the block is PSD will lead to a rank-deficient matrix. This condition is met whenever the last row $(1/2,...,1/2,1+\alpha):=\mathbf{w}$ is orthogonal to the orthogonal complement of the renaming d1d-1 rows: $(-1/d,...,-1/d,1):=\mathbf{v}$. The equality $\mathbf{v}\cdot\mathbf{w}=0$ is fulfilled by $\alpha_{\min}=-(d+1)/(2d)$. Thus, for $\alpha\geq-(d+1)/(2d)$, the last block is PSD as well.
From the previous considerations, we conclude $\alpha_{-}=-(d+1)/(2d)$ and $\alpha_{+}=1$. Finally, Eq. (27) follows from Lemma 1 with the announced bounds on $\alpha$. ∎
Theorem 10.
Let $\rho_{S}\in\mathcal{S}(\mathbb{C}^{d}\otimes\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ be a symmetric state of a $3$-qudit system with $\{\lambda_{k}\leq\lambda_{k+1}\}_{k=0,\cdots,D_{S}-1}$ its corresponding eigenvalues in non-decreasing order, with $D_{S}=d\cdot(d+1)\cdot(d+2)/6$. If at least one of the following inequalities holds:
$\lambda_{0}\geq\frac{1}{D_{S}+\frac{2}{3}}\quad\text{or}\quad\lambda_{D_{S}-1}\leq\frac{1}{D_{S}-\frac{d+2}{3d}},$ (28)
then $\rho_{S}$ is symmetric absolutely PPT.
Moreover, if
$\frac{3d+2}{d+2}\sum_{i=0}^{c-1}\lambda_{i}+\left[D_{S}-\frac{3d+2}{d+2}\cdot c+\frac{2}{3}\right]\cdot\lambda_{c}\geq 0$ (29)
where $c=\left\lceil\frac{3\cdot D_{S}\cdot(2+d)-7d-2}{3\cdot(3d+2)}\right\rceil$, then $\rho_{S}$ is symmetric absolutely PPT.